### 18 Mar 2026 v1

# Zipper-LoRA: Dynamic Parameter Decoupling for Speech-LLM based Multilingual Speech Recognition

Abstract
Speech Large Language Models (Speech-LLMs) have emerged as a powerful approach for automatic speech recognition (ASR) by aligning speech encoders with large language models. However, adapting these systems to multilingual settings with imbalanced data distributions remains challenging. In such scenarios, a stability-plasticity dilemma often arises: fully shared Parameter-Efficient Fine-Tuning (PEFT) can cause negative inter-lingual interference for under-represented languages, while fully language-specific tuning limits the cross-lingual beneficial knowledge transfer needed for low-resource tasks. To address this, we propose Zipper-LoRA, a novel rank-level decoupling framework with three variants (Static, Hard, and Soft) that dynamically synthesizes LoRA updates from shared and language-specific subspaces. By using a lightweight language-conditioned router, Zipper-LoRA dynamically controls the contribution of each subspace at the LoRA rank-level, enabling fine-grained sharing where languages are compatible and strict decoupling when conflicts occur. To further stabilize optimization under imbalanced data, we propose a two-stage training strategy with an Initial-B warm-start that significantly accelerates convergence. Experiments on a 12-language mixed-resource setting show that Zipper-LoRA consistently outperforms both fully shared and independent baselines, particularly in extremely low-resource scenarios. Moreover, we demonstrate that these gains are robust across both chunked and non-chunked encoder configurations, confirming the framework’s reliability for practical, large-scale multilingual ASR. Our code and data will be available at https://github.com/YuCeong-May/Zipper-LoRA for reproducibility.
I Introduction
The rapid progress of Large Language Models (LLMs) has fundamentally reshaped the landscape of modern artificial intelligence. Recent foundation models [40, 1, 50] demonstrate emergent reasoning and generation capabilities, motivating a paradigm shift towards architectures that bridge powerful speech encoders with LLM backbones via projection interfaces [2, 4, 49, 42, 7]. These unified Speech-LLM systems [44, 48, 17] not only achieve high-performance Automatic Speech Recognition (ASR) but also enable complex instruction-following behaviors for speech-centric reasoning tasks.
Despite these advancements, robust and fast domain adaptation of a higher-resource well-trained Speech-LLM model to low-resource or resource-imbalanced multilingual ASR scenarios remains challenging. Current Speech-LLMs are predominantly optimized for high-resource languages (e.g., English, Mandarin), when extended to a long-tailed data distribution that containing numerous low-resource languages, ASR performance often degrades significantly [45]. A primary bottleneck is inter-lingual interference: gradient updates dominated by high-resource training data may distort the shared representations required for under-represented languages. This issue essentially reflects the classic stability-plasticity dilemma in continual learning and multi-task adaptation [34]: an ideal model needs to stay stable enough to handle shared acoustic features while remaining flexible enough to pick up on specific phonological and acoustic differences between languages.
For an ASR adaptation task, directly fine-tuning billion-parameter Speech-LLMs is unrealistic due to the massive computation and memory required. This is why Parameter-Efficient Fine-Tuning (PEFT) has become the standard, with Low-Rank Adaptation (LoRA) [16] being the most common choice. However, using PEFT in diverse multilingual environments brings out a fundamental conflict between reducing inter-lingual interference and encouraging cross-lingual positive knowledge transfer. A typical design shares a single LoRA module across all languages [31], but this fully shared parameterization entangles optimization trajectories; when training data is imbalanced, high-resource languages dominate the shared space and cause negative knowledge transfer [55]. On the other hand, language-specific LoRA assigns an independent module per language [56] eliminates inter-lingual interference but blocks cross-lingual positive transfer. Low-resource languages, lacking sufficient supervision, fail to benefit from universal acoustic and linguistic features (e.g., shared phonemes or articulatory patterns) learned from data-rich languages, resulting in suboptimal generalization.
Recent research works have explored dynamic and modular LoRA variants to bridge this gap. Such as, the proposed Mixture-of-Experts (MoE) adaptions [21, 30, 20] attempt to scale model capacity via multiple experts and routing mechanisms. However, most MoE approaches operate at a coarse granularity (e.g., selecting entire adapter modules or layers) and typically rely on token-level routing that does not explicitly guarantee the decoupling of linguistic attributes. Other methods, such as FlyLoRA [57] or DyLoRA [43], focus on adjusting effective rank for compression or efficiency, yet they lack a mechanism to explicitly distribute capacity between shared and specific components based on language identity. The key research challenge lies in designing a fine-grained mechanism that maximizes the sharing of transferable cross-lingual acoustic knowledge while strictly isolating harmful inter-lingual interference.
Figure 1: Overall Speech-LLM backbone consisting of a speech encoder, a modality projector, and a decoder-only LLM.
To address this challenge, we propose Zipper-LoRA, a novel rank-level dynamic decoupling framework for multilingual Speech-LLM system adaptation. Drawing inspiration from the mechanical action of a zipper, our method dynamically synthesizes the LoRA adaptation matrix by interlocking two complementary subspaces: 1) a Shared Subspace that captures universal acoustic and linguistic regularities, and 2) a Specific Subspace that models language-specific phonological patterns and acoustic atrributes. Unlike static MoE, Zipper-LoRA employs a lightweight router conditioned on language identity to control the contribution of these components at the fine-grained rank-level. By dynamically ‘zipping’ together shared features and ‘unzipping’ conflicting ones at the rank level, Zipper-LoRA provides a flexible solution to balance stability and adaptability. This fine-grained control allows the model to pool cross-lingual knowledge where languages align while blocking inter-lingual interference where they differ.
We evaluate Zipper-LoRA on a standard, high-resource well-trained Speech-LLM architecture that connects a speech encoder, a trainable projector and a text LLM, and adapt it to a 12-language mixed-resource multilingual ASR setting. Our extensive results demonstrate that Zipper-LoRA delivers substantial performance gains. It improves significantly over fully shared LoRA by alleviating inter-lingual interference, and compared with fully independent-LoRA, it achieves better performance while using fewer parameters, confirming its ability to unlock positive transfer with higher efficiency. Our primary contributions are summarized as follows:
  • Dynamic Rank-Level LoRA Decoupling Framework: We propose Zipper-LoRA, a novel PEFT architecture that dynamically synthesizes LoRA adaptation matrices by “zipping” together shared and language-specific subspaces at the rank level. We introduce three distinct variants: Zipper-LoRA-Static (hard partitioning), Zipper-LoRA-Hard (binary column selection), and Zipper-LoRA-Soft (dynamic rank-wise mixing), to provide a flexible range of solutions. This fine-grained composition allows the model to better balance cross-lingual knowledge sharing with the isolation of inter-lingual interference, significantly outperforming coarse-grained MoE approaches.
  • LID-Aware Contextual Routing: We propose a lightweight, rank-level routing mechanism powered by Whisper-derived language identity (LID) embeddings. Unlike traditional MoE methods that switch between entire modules, our router performs fine-grained composition by dynamically controlling the contribution of shared and language-specific subspaces at the individual rank level, ensuring optimal knowledge transfer tailored to each language’s needs.
  • Two-Stage Training with Initial-B Warm-start: To stabilize optimization under imbalanced multilingual data, we develop a robust training strategy. By decoupling cross-modal alignment from language-specific adaptation and employing an Initial-B warm-start, which initializes the low-rank up-projection from a converged dynamic solution, we significantly accelerate convergence and ensure more stable learning of both shared and specific subspaces.
  • Robustness for Practical Deployment: We verify that Zipper-LoRA’s performance gains are consistent across diverse encoder configurations, including both chunked and non-chunked processing. By demonstrating that the framework remains effective regardless of input processing constraints, we confirm its reliability for practical, large-scale multilingual ASR applications.
II Related Work
II-A Large Language Models for Speech Processing
Speech processing has increasingly shifted from task-specific pipelines toward unified Speech Large Language Models (Speech-LLMs), where a speech/audio encoder interfaces with a text LLM through a lightweight alignment module (e.g., linear/MLP projection or a small cross-modal adapter). Earlier systems often followed a cascade paradigm that combines an ASR front-end with a text-only LLM for downstream reasoning, while recent end-to-end Speech-LLMs integrate speech perception and language generation more tightly and support instruction-following behaviors beyond speech recognition [53, 39, 8, 42]. A common design builds upon strong pretrained speech encoders trained on large-scale multilingual speech-text data (e.g., Whisper) [37], and connects them to an LLM backbone via a lightweight interface for representation alignment and conditioning [22, 28]. Along this line, speech-augmented LLMs such as SALM [8] and SALMONN [42] couple audio encoders with (often frozen) LLM backbones and introduce lightweight adaptation modules (e.g., LoRA) to enable ASR and broader audio-language understanding. Representative audio-language foundation models include Qwen-Audio [9] and speech- centric variants such as WavLLM [18], which further demonstrate the feasibility of unifying speech perception with LLM-style generation. Despite strong general capabilities, these models are often biased toward high-resource languages that dominate the pretraining distribution, and adapting Speech-LLMs to multilingual settings with long-tailed language distributions remains challenging, especially for low-resource languages with limited supervision [32, 35, 5].
II-B Parameter-Efficient Fine-Tuning in Multilingual ASR
Fine-tuning billion-parameter models is computationally expensive and can be unstable in data- limited regimes. Parameter-efficient fine-tuning (PEFT) therefore plays a central role in adapting large models to specific domains and languages [14, 15, 26, 19]. Among PEFT methods, low-rank adaptation (LoRA) has become a widely adopted choice due to its simplicity, training stability, and minimal inference overhead [16]. Recent variants further improve practicality and performance, including activation-scaling approaches such as IA3 [27], quantized fine-tuning such as QLoRA [10], and adaptive/weight- decomposed variants such as AdaLoRA and DoRA [54, 29]. In multilingual ASR, LoRA-style adaptation exposes a recurring tension between cross-lingual sharing and language-specific specialization. A shared adaptation module encourages collaborative acoustic and linguistic knowledge transfer by reusing parameters across languages; however, it remains vulnerable to negative interference under imbalanced multilingual training, where updates dominated by high- resource languages can suppress learning signals for low-resource languages [46]. Conversely, language-conditioned or language-specific adaptation reduces such interference but may limit cross-lingual positive transfer and increase the parameter footprint with the number of languages [23, 21, 41]. These observations motivate designs that explicitly allocate capacity to both shared and language-specific components, aiming to retain cross-lingual knowledge transfer while controlling interference in long-tailed multilingual ASR.
II-C Dynamic LoRA Variants and Mixture-of-Experts
To increase the expressivity of parameter-efficient adaptation, recent work has explored dynamic and modular designs. Dynamic-rank approaches such as DyLoRA[43] train LoRA blocks that can be truncated to different effective ranks without expensive search, enabling flexible rank selection. Beyond rank dynamics, Mixture-of-Experts (MoE) style LoRA introduces multiple expert updates and uses routing mechanisms to select or combine experts for each input [47, 25, 51]. While effective for scaling adaptation capacity, many MoE-LoRA variants operate at relatively coarse granularity (e.g., selecting experts at layer/module level), and routing can suffer from imbalance or collapse without careful regularization [12]. Parallel lines of work explore alternative parameterizations of low-rank updates, such as FourierFT that learns a small set of spectral coefficients to reconstruct weight updates efficiently [13], and rank-wise MoE-inspired designs such as FlyLoRA that aim to mitigate interference and improve decoupling [57]. Overall, these methods primarily target efficiency and general expressivity, but they are typically language-agnostic and do not explicitly enforce a boundary between universal acoustic knowledge and language-specific characteristics. In contrast, our work targets multilingual decoupling in LLM-based speech recognition with explicit language awareness and fine-grained control: Zipper-LoRA performs rank-level dynamic routing to allocate LoRA capacity between shared and language-specific subspaces conditioned on language identity, enabling selective sharing of transferable acoustic regularities while isolating language-specific interference.
III Foundations of Speech-LLM and Adaptation Paradigms
In this section, we establish the technical foundations for multilingual Speech-LLM adaptation. We begin by defining the backbone architecture and the language-specific prompting mechanism. Then, we provide a comprehensive overview of representative Parameter-Efficient Fine-Tuning (PEFT) paradigms, ranging from foundational LoRA structures to recent rank-wise Mixture-of-Experts (MoE) adaptations.
III-A Speech-LLM Architecture
Figure 2: Language specific prompts. All these prompts have the same meaning: “Please transcribe the audio content into text.” but are written in specific languages based on the language given for a speech.
Figure 3: Illustration of three representative PEFT frameworks for multilingual ASR Speech-LLM adaptation: Vanilla-LoRA (a), Independent-LoRA (b), and FlyLoRA (c).
Our system adopts a unified Speech-LLM paradigm. As illustrated in Fig. 1, the backbone consists of three primary components: a speech encoder $\mathcal{E}$, a modality projector $\mathcal{P}$, and a large language model $\mathcal{M}$. The process begins with the general formulation of audio-to-text transformation. Given an input speech sequence $\mathbf{X}$, the speech encoder first extracts high-level acoustic representations:
$\mathbf{H}_{enc}=\mathcal{E}(\mathbf{X})$ (1)
To ensure these features are interpretable by the language model, the modality projector $\mathcal{P}$ maps them into the LLM’s token embedding space:
$\mathbf{H}_{proj}=\mathcal{P}(\mathbf{H}_{enc})$ (2)
In a standard setting, the decoder-only LLM $\mathcal{M}$ directly takes these projected embeddings as input to produce the target text sequence $\mathbf{Y}$ through $\mathbf{Y}=\mathcal{M}(\mathbf{H}_{proj})$, optimized via the standard cross-entropy loss.
However, to effectively handle multilingual ASR discrimination, we enhance this standard pipeline by introducing a Language-Specific Prompting mechanism. As shown in Fig. 2, we construct a set of prompts that share the same semantic meaning, namely “Please transcribe the audio content into text.” These prompts are expressed in different languages according to the target speech language. By providing these explicit linguistic cues, we enable the LLM to generate transcriptions in the intended language, thereby reducing language-switching errors. Formally, the prompt is introduced in the text embedding space before the projected speech representations. Let $\mathrm{Emb}(\cdot)$ denote the LLM token embedding lookup and $\oplus$ denote sequence concatenation. For a target language $l$, we first construct the prompt token sequence $\mathbf{T}_{prompt}^{(l)}$ and obtain its embedding representation:
$\mathbf{E}_{prompt}^{(l)}=\mathrm{Emb}\!\left(\mathbf{T}_{prompt}^{(l)}\right)$ (3)
The prompt embeddings are then concatenated with the projected speech representations $\mathbf{H}_{proj}$ to form the final input sequence to the LLM:
$\mathbf{Z}^{(l)}=\mathbf{E}_{prompt}^{(l)}\oplus\mathbf{H}_{proj}.$ (4)
Conditioned on this combined representation, the decoder-only LLM $\mathcal{M}$ generates the target transcription as,
$\mathbf{Y}=\mathcal{M}(\mathbf{Z}^{(l)})$ (5)
III-B Representative PEFT Adaptations of Speech-LLM
To adapt the large-scale Speech-LLM architecture detailed in Sec. III-A for multilingual ASR, full parameter fine-tuning is often computationally expensive and memory intensive. Parameter-efficient fine-tuning (PEFT) methods therefore provide a practical alternative by updating only a small set of additional parameters while keeping the pretrained backbone frozen. Among various PEFT techniques, LoRA-based adaptation has emerged as one of the most widely used approaches for both LLMs and multimodal systems. Building upon this framework, we employ a dual-side LoRA strategy, applying LoRA variants to the speech encoder and standard LoRA on the LLM, to enable efficient multilingual adaptation. As shown in Fig.3, we review three representative PEFT frameworks, ranging from foundational approaches to recent advancements, which serve as strong baselines and the direct motivation for our proposed method.
Technically, for a frozen weight matrix $W_{0}\in\mathbb{R}^{d_{out}\times d_{in}}$, LoRA adapts it by introducing a low-rank update:
$W=W_{0}+\Delta W$ (6)
where the layer output is computed as $W_{0}\mathbf{x}+\Delta W\mathbf{x}$. Different PEFT variants mainly differ in how $\Delta W$ is parameterized and shared across languages.
Vanilla-LoRA: As the most foundational approach, Vanilla-LoRA applies a single low-rank adapter that is fully shared across all languages. As shown in Fig. 3(a), the LoRA branch parameterizes a rank-$r$ update as:
$\Delta W=\frac{\alpha}{r}BA$ (7)
where $A\in\mathbb{R}^{r\times d_{in}}$ and $B\in\mathbb{R}^{d_{out}\times r}$ are shared low-rank factors (down-/up-projection), and $\alpha$ is a scaling hyper-parameter. Since both $A$ and $B$ are fully shared across languages, the same low-rank subspace is used to adapt $W_{0}$. While this promotes maximum cross-lingual knowledge transfer, it is highly susceptible to negative inter-lingual interference. Under imbalanced multilingual training, the shared subspace tends to be dominated by high-resource languages, which can suppress learning signals for low-resource languages [46].
Independent-LoRA: To eliminate such inter-lingual interference, Independent-LoRA represents a traditional design choice that enforces strict parameter isolation. As illustrated in Fig. 3(b), while the pretrained $W_{0}$ remains shared and frozen, a dedicated LoRA module is assigned to each language $l\in\{1,\dots,L\}$:
$\Delta W^{(l)}=\frac{\alpha}{r}B^{(l)}A^{(l)}$ (8)
This design effectively reduces inter-lingual interference by preventing parameter sharing in the LoRA subspace, but it also weakens positive knowledge transfer across languages, which can hurt low-resource languages.
FlyLoRA: A more recent advancement, FlyLoRA [57], introduces a rank-wise mixture-of-experts (MoE) mechanism into the adaptation process. As shown in Fig. 3(c), it treats each rank component as an individual expert and dynamically activates a sparse subset of ranks via an implicit routing mechanism. Specifically, it employs a sparse and frozen down-projection matrix $A$ to compute routing scores $\mathbf{y}=A\mathbf{x}+\mathbf{d}$, where $\mathbf{d}\in\mathbb{R}^{r}$ is a learnable bias term for load balancing. Based on these scores, a top-$k$ operator selects the active indices $\mathcal{I}_{\text{top}k}\subseteq\{1,\dots,r\}$. The adapter weight is then constructed by aggregating the top-$k$ active components:
$\Delta W=\frac{\alpha}{r}\sum_{i\in\mathcal{I}_{\text{top}k}}\mathbf{b}_{i}\mathbf{a}_{i}^{\top}$ (9)
where $\mathbf{b}_{i}$ and $\mathbf{a}_{i}^{\top}$ are the $i$-th column and row of the up- and down-projection matrices, respectively. Although FlyLoRA was originally proposed for general task decoupling in LLMs, its unique rank-wise MoE structure serves as a strong baseline and a direct inspiration for our work. However, since FlyLoRA relies on implicit routing without explicit language cues, it lacks a dedicated mechanism to decouple language-specific specialization from shared acoustic knowledge. This limitation motivates our proposed Zipper-LoRA, which explicitly harmonizes these two components.
IV Proposed Zipper-LoRA
Figure 4: Overview of the proposed Zipper-LoRA. A language-aware router outputs rank-wise mixing weights from language embeddings to construct $B_{\text{merged}}^{(l)}$ for multilingual ASR adaptation.
As discussed in the above sections, the adaptation of multilingual ASR Speech-LLMs poses two competing requirements: 1) The model needs shared capacity to facilitate cross-lingual knowledge transfer, which is important for improving low-resource language performance; 2) It requires language-discriminative capacity to mitigate inter-lingual interference, particularly under the imbalanced multilingual data training. Throughout this process, the adaptation must remain parameter-efficient to stay scalable for large-scale frozen backbones, i.e., introducing only a small number of trainable parameters on top of a frozen backbone.
To harmonize these objectives, we propose Zipper-LoRA, a family of PEFT modules designed to explicitly allocate and integrate shared and language-specific rank components in a language-aware manner. As illustrated in Fig. 4, all Zipper-LoRA variants leverage a globally shared down-projection matrix $A$, ensuring common acoustic-linguistic feature extraction. The fundamental distinction between these variants lies in the rank-wise construction of the up-projection matrix $B$. Intuitively, Zipper-LoRA “zips” together shared and language-specific columns of $B$ to form an effective rank-$r$ adapter for each language. This integration can be realized through either a static allocation or a dynamic rank-wise composition.
To provide a comprehensive understanding of the proposed framework, in the following sections, we first establish the definition of the shared down-projection matrix $A$ and the rank-wise construction of $B$. We then present the proposed Zipper-LoRA family, ranging from the static language-wise hard split to dynamic binary selection and soft mixing mechanisms. This is followed by the LID-aware contextual routing that manages these rank-wise compositions, concluding with the specialized training strategies tailored for optimizing Zipper-LoRA, such as segment-wise encoding and parameter initialization.
IV-A Definition: Shared $A$ and Rank-wise Construction of $B$
Let $A\in\mathbb{R}^{r\times d_{in}}$ be the shared down-projection matrix. Zipper-LoRA constructs an effective up-projection matrix $B_{\text{merged}}\in\mathbb{R}^{d_{out}\times r}$ by combining a shared bank $B_{\text{shared}}$ and a language-specific bank $B_{\text{spec}}^{(l)}$. The resulting adaptation is applied as
$\Delta W^{(l)}\;=\;\frac{\alpha}{r}\,B_{\text{merged}}^{(l)}\,A$ (10)
where $\Delta W^{(l)}$ denotes the LoRA-induced weight update for language $l$, and $B_{\text{merged}}^{(l)}$ is instantiated differently by the three variants below. In all cases, $B_{\text{merged}}$ is formed rank-wise, which matches the visualization in Fig. 4(a)-(c).
IV-B Zipper-LoRA-Static (Language-wise Hard Split)
Zipper-LoRA-Static implements a simple MoE-style specialization by hard partitioning rank components into shared and language-specific parts. As shown in Fig. 4(a), we assume the target rank is $r$ (e.g., r=32r{=}32), and split it into $r_{s}$ shared ranks and $r_{p}$ language-specific ranks, with $r_{s}+r_{p}=r$ (e.g., $r_{s}{=}16$, $r_{p}{=}16$). We learn a shared matrix $B_{\text{shared}}\in\mathbb{R}^{d_{out}\times r_{s}}$ and a language-specific matrix $B_{\text{spec}}^{(l)}\in\mathbb{R}^{d_{out}\times r_{p}}$ for each language $l$. The merged up-projection is formed by concatenation:
$B_{\text{merged}}^{(l)}\;=\;\big[\,B_{\text{shared}},B_{\text{spec}}^{(l)}\,\big]\in\mathbb{R}^{d_{out}\times r}$ (11)
This design preserves a fixed amount of shared capacity while allocating a dedicated subspace for each language.
IV-C Zipper-LoRA-Hard (Dynamic Binary Column Selection)
Zipper-LoRA-Hard enables language-dependent specialization by selecting rank components from the language-specific bank in a binary manner. As shown in Fig. 4(b), we learn a shared bank $B_{\text{shared}}\in\mathbb{R}^{d_{out}\times r}$ and a language-specific bank $B_{\text{spec}}^{(l)}\in\mathbb{R}^{d_{out}\times r}$ (both of rank $r$). Given a target language $l$, a language-aware router takes the corresponding language embedding $\mathbf{e}^{(l)}$ as input and outputs a rank-wise gating vector $\mathbf{p}^{(l)}\in[0,1]^{r}$ . We then obtain a binary selection mask by thresholding:
$s_{i}^{(l)}=\mathbb{I}\left[p_{i}^{(l)}\geq\tau\right],\qquad\mathbf{s}^{(l)}\in\{0,1\}^{r}$ (12)
where $\tau$ is a fixed threshold. The merged up-projection is constructed by taking the selected columns from $B_{\text{spec}}^{(l)}$ and the remaining columns from $B_{\text{shared}}$:
$B_{\text{merged}}^{(l)}\;=\;\mathrm{Zip}\Big(B_{\text{shared}},\,B_{\text{spec}}^{(l)},\,\mathbf{s}^{(l)}\Big)$ (13)
where $\mathrm{Zip}(\cdot)$ denotes the rank-wise “zipper” operation shown in Fig. 4(b): the $i$-th column of $B_{\text{merged}}^{(l)}$ is taken from $B_{\text{spec}}^{(l)}$ if $s_{i}^{(l)}{=}0$, otherwise from $B_{\text{shared}}$. This yields an effective rank-$r$ adapter whose shared vs. language-specific columns are controlled by the binary mask $\mathbf{s}^{(l)}$. During gradient backpropagation, we adopt the straight-through estimator (STE) [6] to approximate gradients through the non-differentiable thresholding operation in Eq.(12).
IV-D Zipper-LoRA-Soft (Dynamic Rank-wise Column Mixing)
Zipper-LoRA-Soft replaces binary selection with continuous rank-wise mixing between shared and language-specific columns, yielding smoother adaptation. As shown in Fig. 4(c), while maintaining the same rank-$r$ bank structure as Zipper-LoRA-Hard, we replace the binary mask $\mathbf{s}^{(l)}$ with a continuous mixing vector $\mathbf{p}^{(l)}\in[0,1]^{r}$ generated by the language-aware router. We interpret $\mathbf{p}^{(l)}$ as the rank-wise proportion assigned to the language-specific columns, and construct the merged up-projection via a weighted summation:
$B_{\text{merged}}^{(l)}\;=\;B_{\text{shared}}\,\mathrm{diag}\!\big(\mathbf{1}-\mathbf{p}^{(l)}\big)\;+\;B_{\text{spec}}^{(l)}\,\mathrm{diag}\!\big(\mathbf{p}^{(l)}\big)$ (14)
where $\mathrm{diag}(\mathbf{p})$ denotes the diagonal matrix with $\mathbf{p}$ on its diagonal, i.e., it performs rank-wise (column-wise) scaling. Equivalently, for each rank index $i\in\{1,\dots,r\}$,
$\mathbf{b}_{\text{merged},i}^{(l)}\;=\;\big(1-p_{i}^{(l)}\big)\,\mathbf{b}_{\text{shared},i}\;+\;p_{i}^{(l)}\,\mathbf{b}_{\text{spec},i}^{(l)}$ (15)
Compared to Zipper-LoRA-Hard, Zipper-LoRA-Soft provides a continuous trade-off between shared and language-specific capacity without requiring discrete rank selection. Here “dynamic” refers to rank-wise adaptive weighting: different rank components can be assigned different mixing weights for each language.
IV-E LID-Aware Contextual Routing
Whisper LID Embedding: In Fig. 4, both the Zipper-LoRA-Hard and Zipper-LoRA-Soft require a language-aware router for the LoRA rank selection, we obtain the language identity (LID) embedding from the Whisper decoder [37]. Specifically, Whisper produces a fixed-dimensional LID representation $\mathbf{e}^{(l)}\in\mathbb{R}^{d_{\text{lid}}}$ for target language $l$, which serves as a compact language-conditioned signal for routing.
Router Architecture: As illustrated by the router block in Fig. 4, we adopt a lightweight router that maps the Whisper LID embedding to rank-wise mixing weights. Specifically, we first normalize the LID embedding and then project it to the LoRA rank dimension:
$\tilde{\mathbf{e}}^{(l)}=\mathrm{LayerNorm}\!\left(\mathbf{e}^{(l)}\right),$ (16)
$\hat{\mathbf{e}}^{(l)}=W_{r}\,\tilde{\mathbf{e}}^{(l)}+\mathbf{b}_{r},\quad\hat{\mathbf{e}}^{(l)}\in\mathbb{R}^{r}$ (17)
where $W_{r}\in\mathbb{R}^{r\times d_{\text{lid}}}$ and $\mathbf{b}_{r}\in\mathbb{R}^{r}$ are trainable parameters. Then, we apply an element-wise sigmoid to obtain rank-wise mixing weights:
$\mathbf{p}^{(l)}=\sigma\!\left(\hat{\mathbf{e}}^{(l)}\right),\quad\mathbf{p}^{(l)}\in[0,1]^{r}$ (18)
The resulting $\mathbf{p}^{(l)}$ is the router output shown in Fig. 4(b) and (c), and it directly controls rank-wise composition between the shared and language-specific banks. In particular, Zipper-LoRA-Soft uses $\mathbf{p}^{(l)}$ as continuous mixing weights in Eq. (14), while Zipper-LoRA-Hard derives a binary mask $\mathbf{s}^{(l)}\in\{0,1\}^{r}$ by thresholding $\mathbf{p}^{(l)}$ (Eq. (12)) and performs rank-wise selection via Eq. (13).
IV-F Training Strategy of Speech-LLM with Zipper-LoRA
This section describes how the model is trained for multilingual ASR adaptation and how the proposed initialization and efficiency strategies are applied. Stage 1 serves as a foundation baseline to obtain a well-calibrated Speech-LLM backbone. All subsequent experiments, including different LoRA variants, are conducted by inserting LoRA modules into the speech encoder and training them in a parameter-efficient supervised fine-tuning (SFT) regime.
Stage 1: Foundation Baseline: In Stage 1, we train an alignment baseline to improve the coupling between the speech encoder and the LLM. Specifically, as the pipeline used in [31], we update the speech encoder parameters and the modality projector, and train LoRA parameters injected into the LLM, while keeping the remaining LLM backbone frozen. This stage provides a strong starting point with improved cross-modal alignment, and is used as the backbone for Stage 2. Notably, Stage 1 does not involve encoder-side LoRA; encoder LoRA is only introduced in Stage 2 for all PEFT variants.
Stage 2: Parameter-Efficient Multilingual ASR SFT: In Stage 2, we freeze the speech encoder weights obtained from Stage 1 and insert LoRA modules into the encoder. We then perform multilingual SFT by updating only the encoder-side LoRA parameters (with different variants), the projector, and the LLM-side LoRA parameters. Under this setting, different PEFT frameworks differ only in how the encoder-side low-rank update is structured and shared across languages, while the rest of the training pipeline remains identical.
Segment-wise (Chunked) Encoder Processing for Training Efficiency: Training on long-form utterances with full-context encoding can be memory and compute-intensive. To improve training throughput, we adopt a segment-wise (chunked) encoder setting. Given an utterance, we split the input audio into fixed-length segments and encode each segment independently using the same encoder. Segment representations are concatenated in temporal order and fed into the projector. This strategy reduces peak GPU memory usage while keeping the Speech-LLM architecture unchanged. In experiments, we report results under both full-context and segment-wise encoder regimes to verify robustness to encoder input processing.
Initial-B Warm-start Initialization: To stabilize optimization of rank-wise routing under long-tailed multilingual supervision, we propose an Initial-B warm-start strategy. In standard LoRA, the up-projection matrix $B$ is typically initialized to zeros, which can delay effective adaptation and make joint learning of routing and multilingual specialization less stable. Instead of zero initialization, we warm-start the $B$ parameters using a converged dynamic Zipper solution.
Specifically, we first train a dynamic Zipper variant with rank-wise mixing under the Stage 2 SFT setting to obtain a converged solution. We then reuse its learned up-projection parameters to initialize subsequent runs: both the shared bank $B_{\text{shared}}$ and the language-specific bank $B_{\text{spec}}^{(l)}$ are initialized from the pretrained dynamic Zipper parameters (and the router can optionally be warm-started as well). This warm-start provides a well-shaped low-rank subspace at the beginning of training, enabling the model to immediately leverage meaningful rank components and improving convergence and training stability for the target setting.
V EXPERIMENTAL CONFIGURATIONS
V-A Datasets
TABLE I: Dataset statistics (duration in hours).
Setting Dataset Language Train(h) Test(h)
Baseline (source) MSR86k[24] English (en) 3000 14.72
French (fr) 3000 12.60
Thai (th) 3000 8.37
WenetSpeech (WS)[52] Chinese (zh) 3000 15.18 $\mid$ 23.06
SFT (target) MLS[36] German (de) 1000 14.29
Spanish (es) 1000 10.00
French (fr) 1000 10.07
MSR86k[24] Russian (ru) 1000 8.16
Vietnamese (vi) 1000 7.25
Italian (it) 500 8.10
LibriSpeech (LS)[33] English(en) 960 5.40 $\mid$ 5.34
Common Voice (CV)[3] Thai (th) 865 4.00
Arabic (ar) 1 4.00
Japanese (ja) 1 4.00
Korean (ko) 1 0.98
Portuguese (pt) 1 1.00
[…]
Table I summarizes the data used in our two-stage training. The foundation baseline stage uses high-resource corpora to build a multilingual ASR initialization, while the SFT stage (stage 2) adapts the model to a 12-language setting with substantial training resource imbalance.
As shown in Table I (Baseline block), we construct the foundation baseline training set with four high-resource languages. We use MSR86k to provide English/French/Thai and WenetSpeech to provide Chinese, and for each language we randomly sample 3000 hours to control the training budget. From the SFT block, we perform multilingual ASR SFT adaptation on 12 target languages with a long-tailed distribution. The higher-resource portion consists of MLS German/Spanish/French/Russian (1000 hours each), MSR86k Vietnamese (1000 hours) and Italian (500 hours), LibriSpeech English (960 hours), and Common Voice Thai (865 hours), where the hours are obtained by random sampling as listed in the table. The lowest-resource portion comes from Common Voice Arabic/Japanese/Korean/Portuguese with 1 hour per language. Some languages overlap across stages but come from different corpora (French, Thai, and English), introducing domain mismatch that can increase cross-lingual interference under long-tailed SFT adaptation.
For evaluation, we follow the official splits of each dataset whenever available. MSR86k provides an official dev split (and no official test split), so we evaluate on its dev set. For MLS, LibriSpeech, and WenetSpeech, we evaluate on the official test splits. For Common Voice, we use randomly sampled evaluation subsets to control the evaluation scale: Arabic and Thai use 4-hour subsets randomly sampled from the official test split, Korean uses the official test split (0.98 hours), and Portuguese uses a 1- hour subset randomly sampled from the official training split as a held-out evaluation set due to the limited data size and the absence of an official test split. All random sampling procedures are fixed and shared across all methods to ensure fair comparisons, and we have released the exact utterance lists for reproducibility.
V-B Configurations
We use Whisper Large-v3 [37] as the speech encoder and Qwen3-1.7B [50] as the decoder-only language model. The modality projector adopts a gated projection module that performs temporal downsampling by a factor of $4$. Specifically, it applies two 1D convolution layers (kernel size $4$, stride $4$) to produce a gate branch and an up-projection branch, fuses them via a SiLU-gated interaction [11], and then applies a linear transformation followed by a residual connection with layer normalization. A final linear projection outputs the speech-conditioned embedding sequence. Combined with the initial $2\times$ downsampling in the speech encoder, this results in an overall $8\times$ temporal reduction, yielding a highly compressed representation at $12.5$ Hz that is fed into the language model.
For parameter-efficient adaptation, we insert low-rank adaptation modules into all linear layers on both the encoder and language-model sides, including the attention projections and feed-forward network layers. We set the rank to $32$ and the scaling factor to $64$ for all methods. For FlyLoRA [57], we activate $8$ rank components. For the chunked encoder regime, we randomly sample chunk durations from $\{1,2,4,8\}$ seconds during training and use a fixed chunk duration of $8$ seconds during decoding; no overlap is used. The non-chunked regime encodes each utterance with full context. All experiments are trained on $4$ NVIDIA A800 GPUs (80 GB) with DeepSpeed ZeRO stage $2$ [38], using a learning rate of $2\times 10^{-5}$, a warmup ratio of $10\%$, and a cosine learning-rate schedule. To improve memory efficiency, we pack training samples by constraining the total sequence length to at most 2048 tokens, where the packed length includes the downsampled speech tokens, the text tokens, and reserved prompt tokens; packing is performed within the same language. After packing, we use a batch size of $16$ for chunked training and a batch size of $10$ for non-chunked training.
For evaluation, we compute character error rate (CER) for Chinese, Japanese, Korean, and Thai, and word error rate (WER) for all other languages, with Whisper text normalization. All experiments in this work follow these configurations unless otherwise specified.
VI Results and Discussions
VI-A Source Domain Results
TABLE II: Foundation baseline WER/CER (%) comparison under chunked vs. non-chunked settings.
MSR86k WenetSpeech
en fr th Meeting Net
Chunked 4.65 9.24 4.20 16.55 9.60
Non-chunked 4.59 9.29 4.05 17.16 10.22
TABLE III: Source-domain performance after SFT on MSR86k (en/fr/th) and Wenetspeech (Meeting/Net) under chunked and non-chunked encoder settings.
MSR86k WenetSpeech
Method en fr th Meeting Net
Chunked
Vanilla-LoRA 4.88 10.74 13.58 18.51 12.12
FlyLoRA 4.87 10.65 7.99 17.57 11.86
Independent-LoRA 4.86 11.07 7.72 17.24 11.63
Zipper-LoRA-Static 4.91 11.35 7.85 18.12 12.11
Zipper-LoRA-Hard 4.79 11.06 7.59 17.49 11.61
Zipper-LoRA-Soft 4.88 11.26 7.79 17.90 11.93
     + initial-B 5.06 11.53 8.05 17.96 12.02
Non-chunked
Vanilla-LoRA 4.40 10.10 12.04 18.40 12.38
FlyLoRA 4.43 10.10 7.10 18.25 12.31
Independent-LoRA 4.59 10.86 7.50 17.58 11.76
Zipper-LoRA-Static 4.64 11.07 7.70 17.68 11.97
Zipper-LoRA-Hard 4.53 10.78 7.40 18.06 11.79
Zipper-LoRA-Soft 4.59 11.02 7.52 17.99 11.90
     + initial-B 4.69 11.24 7.69 17.83 11.94
Note: “+ initial-B” denotes warm-start for Zipper-LoRA-Soft.
Table II reports pretrained-stage foundation baseline ASR performance under chunked and non-chunked encoder settings. On MSR86k (English, French, and Thai), the two settings are essentially on par, with only minor language-dependent differences (non-chunked slightly favors English/Thai, while chunked is marginally better on French). In contrast, on the WeNetSpeech Meeting and Net sets, the chunked encoder consistently yields lower error rates, suggesting better robustness in longer-form and more heterogeneous conditions. Overall, these results indicate that chunking does not introduce systematic degradation on the source domain, while offering a small but consistent advantage on more challenging evaluation scenarios.
Table III presents the source-domain ASR performance after SFT adaptation under both chunked and non-chunked encoder configurations. A comparison between the foundation baseline (Table II) and the post-SFT results reveals a general performance degradation across all evaluated languages. This decline is primarily caused by domain mismatch and cross-lingual interference during the second-stage adaptation. Although languages such as English, French, and Thai are present in both stages, the SFT stage uses different corpora (e.g., LibriSpeech and Common Voice) compared to the foundation stage (MSR86k). This introduces significant distributional shifts; for instance, in the Vanilla-LoRA (Chunked) setting, the error rate for Thai increases sharply from 4.20% to 13.58%. Such results suggest that the long-tailed distribution of the 12-language SFT task leads to severe interference, resulting in “catastrophic forgetting” of the high-resource knowledge learned during the foundation stage.
The degree of forgetting varies significantly across different PEFT strategies. Vanilla-LoRA shows the most obvious performance drop, particularly on the Thai and WenetSpeech tasks, indicating that a shared low-rank space is not enough to protect source-domain knowledge from the noise of multi-target adaptation. While Independent-LoRA and FlyLoRA reduce this forgetting by decoupling parameters with Independent-LoRA showing good stability on the WenetSpeech Meeting and Net sets, they often lack the flexibility to maintain optimal performance across all source tasks at the same time.
Our proposed Zipper-LoRA framework, especially the Zipper-LoRA-Hard variant, shows better performance in managing the trade-off between target-domain adaptation and source-domain preservation. In the chunked encoder setting, Zipper-LoRA-Hard achieves the lowest error rates for English (4.79%), Thai (7.59%), and WenetSpeech Net (11.61%) among all compared SFT methods. This performance gain highlights the effectiveness of the Zipper-LoRA routing mechanism in isolating interference from the long-tailed target distribution. Furthermore, the way Zipper-LoRA and chunked encoding work together is clear in long-form scenarios like WenetSpeech, where it consistently performs better than non-chunked alternatives. These results confirm that Zipper-LoRA-Hard provides a strong protection, effectively “zipping” the foundation and adaptation layers to reduce the negative impact of domain mismatch.
Figure 5: Performance comparison of Zipper-LoRA-Soft (+ initial-B) and other LoRA-based methods on the 12 target languages in the SFT stage under the non-chunked setting, using $(1-\mathrm{WER/CER})\%$ as the metric. Detailed numerical results are provided in Tables V and VI.
VI-B Target Domain Results
Fig.5 compares multilingual ASR adaptation performance on the 12 target languages in the supervised fine-tuning stage under the non-chunked encoder setting, using $(1-\mathrm{WER/CER})\%$ as the score. Overall, Zipper-LoRA-Soft (+initial-B) achieves a consistently larger enclosed area than Vanilla-LoRA, Independent-LoRA and FlyLoRA, indicating better balanced performance across languages. This suggests that rank-level composition with controlled sharing effectively mitigates inter-lingual interference while preserving corss-lingual transferable acoustic representations.
VI-B1 Results on high-resource target languages
TABLE IV: WER/CER(%) on high-resource target languages with chunk-encoder.
Methods CommonVoice LibriSpeech MLS MSR86k
th Clean Other Avg de es fr it ru vi
Vanilla-LoRA 1.47 2.77 5.88 4.32 9.57 6.31 8.31 12.57 9.59 6.64
FlyLoRA 1.57 2.80 5.90 4.35 9.71 6.48 8.28 12.92 9.70 6.75
Independent-LoRA 1.30 2.60 5.63 4.11 9.08 5.67 8.02 11.65 9.40 6.33
Zipper-LoRA-Static 1.35 2.73 5.58 4.15 9.20 6.10 8.15 11.87 9.31 6.41
Zipper-LoRA-Hard 1.35 2.62 5.53 4.07 9.15 5.69 8.03 11.71 9.44 6.36
Zipper-LoRA-Soft 1.31 2.61 5.54 4.07 9.15 5.62 8.03 11.75 9.37 6.32
           + initial-B 1.26 2.56 5.45 4.00 8.93 5.31 7.90 11.10 9.21 6.15
TABLE V: WER/CER(%) on high-resource target languages with non-chunk encoder.
Methods CommonVoice LibriSpeech MLS MSR86k
th Clean Other Avg de es fr it ru vi
Vanilla-LoRA 1.48 2.66 5.65 4.15 8.52 5.27 8.26 11.86 9.05 6.31
FlyLoRA 1.49 2.70 5.68 4.19 9.04 5.79 8.13 12.36 9.23 6.51
Independent-LoRA 1.36 2.53 5.44 3.98 8.36 5.05 8.09 11.66 9.02 6.13
Zipper-LoRA-Static 1.30 2.52 5.45 3.98 8.53 5.00 7.96 11.95 8.97 6.18
Zipper-LoRA-Hard 1.35 2.56 5.49 4.02 8.39 4.86 7.82 11.36 8.94 6.19
Zipper-LoRA-Soft 1.36 2.55 5.47 4.01 8.38 4.86 7.84 11.41 8.87 6.11
           + initail-B 1.32 2.53 5.38 3.96 8.20 4.84 7.50 11.09 8.79 5.94
TABLE VI: WER/CER(%) on low-resource Common Voice languages under chunked vs. non-chunked settings.
Methods Chunk Non-chunk
ar ja ko pt ar ja ko pt
Vanilla-LoRA 62.32 48.93 38.37 33.49 53.08 46.28 36.03 29.30
FlyLoRA 65.20 49.75 39.02 33.58 55.01 46.86 35.99 30.16
Independent-LoRA 45.24 37.84 25.98 25.15 41.01 37.53 24.53 23.82
Zipper-LoRA-Static 42.36 37.51 25.65 23.26 40.32 36.61 23.39 24.61
Zipper-LoRA-Hard 43.07 37.10 25.80 23.86 40.89 36.55 24.61 23.96
Zipper-LoRA-Soft 41.72 35.77 24.37 23.89 40.07 35.66 22.55 23.47
           + initial-B 38.64 35.21 21.87 24.29 36.94 34.04 21.35 23.22
Table IV and Table V report SFT ASR performance on higher-resource target languages under the chunked and non-chunked encoder settings, respectively, including Common Voice (Thai), LibriSpeech (Clean, Other and Average), Multilingual LibriSpeech (German, Spanish, French and Italian) and MSR86k (Russian and Vietnamese). Across both the chunked and non-chunked encoder settings, the Zipper-LoRA family consistently delivers strong performance and exhibits the same overall trend. In particular, Zipper-LoRA-Soft closely matches Independent-LoRA and even surpasses it on several high-resource languages. For example, under the non-chunked setting, Zipper-LoRA-Soft outperforms Independent-LoRA on Spanish, French, Italian, Russian and Vietnamese. Moreover, adopting the +initial-B warm-start strategy further improves performance, yielding a comprehensive advantage over Independent-LoRA. This is also shown by the radar visualization in Fig.5, where Zipper-LoRA variants achieve a larger overall coverage, indicating more balanced gains across languages. By contrast, Vanilla-LoRA and FlyLoRA underperform compared to the proposed Zipper-LoRA, likely due to the inter-lingual interference introduced by parameter sharing.
VI-B2 Results on low-resource target languages
In the extremely low-resource adaptation tasks (only 1 hour of fine-tuning data per language), Zipper-LoRA exhibits a clear advantage over both the fully shared baseline (Vanilla-LoRA) and the fully decoupled alternative (Independent-LoRA). As summarized in Table VI, all Zipper-LoRA variants achieve lower WER/CER on the four lowest-resource Common Voice languages (Arabic, Japanese, Korean and Portuguese), indicating that structured cross-lingual sharing is crucial when supervision is extremely limited. While Independent-LoRA eliminates interference via complete parameter decoupling, it simultaneously removes the ability to exploit acoustic knowledge transferable representations across languages; therefore, its adaptation capacity is constrained under 1-hour fine-tuning. In contrast, Zipper-LoRA preserves beneficial cross-lingual sharing while controlling inter-lingual interference through rank-level composition, enabling low-resource languages to borrow statistical strength from other languages without being overwhelmed. Meanwhile, densely shared approaches such as Vanilla-LoRA and FlyLoRA are more sensitive to negative transfer: with fully shared adaptation parameters, gradients from high-resource languages can dominate and distort the updates for long-tail languages, leading to large performance degradation. Overall, these results suggest that the proposed Zipper-LoRA with fine-grained composition provides a better trade-off between cross-lingual knowledge transfer and inter-lingual interference than either fully shared or fully decoupled designs, and is particularly effective for stabilizing Speech-LLM adaptation in the long-tail ASR adaptation setting.
VI-B3 Scalability from extremely-low to low-resource target languages
Figure 6: Performance gap visualization from extremely-low to low-resource target languages. Bars show relative performance change vs. Vanilla-LoRA; positive values indicate improvement and negative values indicate degradation. The 1-hour setting uses 1 hour per language as in Table VI; for the 30-hour setting, Arabic, Japanese, and Korean are from Common Voice (Korean uses 6.51 hours due to the limited official data size), while Portuguese is from MLS.
Fig. 6 shows the visualized relative performance change of each method compared to Vanilla-LoRA. Positive bars indicate improvements, while negative bars show a drop in performance. This visualization examines whether the benefits of the proposed methods hold when the training data for low-resource target languages increases from 1 hour (extremely low-resource in Table VI) to 30 hours (low-resource).
Across both chunked and non-chunked regimes, Zipper-LoRA variants consistently show gains over Vanilla-LoRA. The most significant improvements are observed in the 1-hour setting, where the lack of data makes structured cross-lingual sharing vital for the model to learn effectively. As the data size increases to 30 hours, the relative improvement gap narrows slightly. This is expected, as the increased supervision allows the model to rely more on language-specific information. However, the Zipper-LoRA-Soft + initial-B variant still maintains a clear advantage over all other methods even at the 30-hour condition. This suggests that our approach is not just a solution for extreme data scarcity, but remains effective as more data becomes available. This is also consistent with the observation obtained in Table V and IV for high-resource target languages ASR adaptation.
In contrast, fully shared adaptations as FlyLoRA can exhibit smaller gains or even degradation, suggesting stronger sensitivity to negative knowledge transfer when low-resource languages are adapted jointly with higher-resource ones. While Independent-LoRA provides strong protection against interference by isolating parameters, it cannot capture the cross-lingual benefits that Zipper-LoRA achieves through its rank-wise composition. Overall, these results further confirm that Zipper-LoRA provides a flexible framework that remains strong across different data scales, effectively managing the transition from extremely low to low-resource conditions.
VI-B4 Robust gains and the effect of encoder context
Comparing Tables IV and V, together with Table VI, shows that the results are robust to the encoder processing strategy. Across all methods, the non-chunked setting consistently yields lower absolute WER than the chunked setting, since chunked attention processes each block independently and prevents the encoder from attending across chunks, thereby limiting the available contextual information. Despite this systematic gap, Zipper-LoRA variants, particularly soft composition and the initial-B warm-start, consistently achieve the lowest or near-lowest error rates across both high-resource benchmarks and the low-resource Common Voice subset. This consistency supports the claim that rank-wise sharing and specialization provide a reliable mechanism for balancing cross-lingual knowledge transfer and inter-lingual interference under different encoder context conditions in practical multilingual adaptation.
VI-C Representation Visualization
Figure 7: Cosine-similarity heatmap of Whisper language-ID (LID) embeddings for the 12 target languages.
We first analyze the language-ID (LID) embeddings provided by Whisper, which encode the model’s prior notion of language similarity and are expected to reflect broad phonetic and acoustic similarities. Fig.7 shows the pairwise cosine-similarity structure among the 12 target languages. We observe distinct high-similarity clusters between linguistically related languages; for example, Japanese (ja) and Korean (ko) exhibit a cosine similarity of 0.41, and Thai (th) and Vietnamese (vi) also reach 0.41. In contrast, more distant language pairs consistently show lower similarity values. This semantic topology indicates that Whisper’s pretrained LID space already organizes languages into meaningful neighborhoods, motivating the use of LID embeddings as a routing anchor: nearby languages are more likely to benefit from knowledge transfer, whereas forcing distant languages to share adaptation capacity increases the risk of inter-lingual interference.
Figure 8: t-SNE visualization of encoder output representations across the 12 target languages.
We then examine the encoder output representations after SFT ASR adaptation. Fig.8 visualizes these representations using t-SNE, where each point corresponds to an utterance embedding and colors indicate languages. A clear contrast emerges between fully shared Vanilla-LoRA and Zipper-LoRA-Soft. Under Vanilla-LoRA (left), multiple languages collapse into partially overlapping or densely entangled regions, indicating that the shared adaptation parameters induce cross-lingual interference; this effect is especially increased for low-resource languages, whose representations are more easily dominated and distorted by updates driven by higher-resource languages, leading to negative knowledge transfer. In contrast, Zipper-LoRA-Soft (right) yields markedly more compact and well-separated language clusters in the embedding space, suggesting that it suppresses harmful intel-lingual interference while preserving language-specific factors. This separation provides qualitative evidence for the effectiveness of rank-wise composition: by enabling fine-grained sharing and specialization, Zipper-LoRA-Soft can exploit transferable structure where beneficial while preventing excessive coupling across unrelated or imbalanced languages, resulting in more stable multilingual adaptation.
VII Conclusion
Multilingual ASR adaptation under long-tailed data distributions must balance cross-lingual transfer and interference. Fully shared parameter-efficient tuning can suffer from negative knowledge transfer, especially for under-represented languages, while fully language-specific tuning reduces intel-lingual interference but sacrifices beneficial sharing and can be severely constrained in the extremely low-resource conditions.
In this work, we introduced Zipper-LoRA, a rank-wise composition framework that decomposes adaptation capacity into shared and language-specific components and combines them through language-conditioned routing. This design enables fine-grained sharing where languages are compatible while preserving specialization to suppress harmful coupling. Across 12 target languages covering high and low-resource conditions, Zipper-LoRA variants consistently achieve strong performance and more balanced gains than shared baselines and fully decoupled alternatives. The improvements are particularly evident on long-tail languages, where controlled sharing provides additional benefit beyond complete parameter decoupling. We further showed that these advantages are robust across different encoder processing settings: although non-chunked encoding yields lower absolute error rates than chunked encoding due to the availability of broader context, the relative ranking and gains of Zipper-LoRA remain consistent. Finally, representation analyses support the proposed mechanism by revealing meaningful language structure in LID embeddings and clearer separation of adapted encoder representations under Zipper-LoRA, consistent with reduced inter-lingual interference.
Overall, our results suggest that rank-wise composition offers an effective and robust approach for practical multilingual adaptation, providing a better trade-off between cross-lingual transfer and inter-lingual interference than either fully shared or fully independent designs. Future work includes scaling to a larger and more diverse set of languages, exploring alternative routing signals beyond LID, and extending the approach to streaming or continual multilingual adaptation settings.
Acknowledgments
This work was supported by the Natural Science Foundation of Shanghai (Grant No. 25ZR1401277) and the National Natural Science Foundation of China (Grant No. 62071302).

### 19 Mar 2026 v2

@@
 where $A\in\mathbb{R}^{r\times d_{in}}$ and $B\in\mathbb{R}^{d_{out}\times r}$ are shared low-rank factors (down-/up-projection), and $\alpha$ is a scaling hyper-parameter.
-Since both $A$ and $B$ are fully shared across languages, the same low-rank subspace is used to adapt $W_{0}$.
+Since both AA and $B$ are fully shared across languages, the same low-rank subspace is used to adapt $W_{0}$.
 While this promotes maximum cross-lingual knowledge transfer, it is highly susceptible to negative inter-lingual interference.
@@
 3(c), it treats each rank component as an individual expert and dynamically activates a sparse subset of ranks via an implicit routing mechanism.
-Specifically, it employs a sparse and frozen down-projection matrix $A$ to compute routing scores $\mathbf{y}=A\mathbf{x}+\mathbf{d}$, where $\mathbf{d}\in\mathbb{R}^{r}$ is a learnable bias term for load balancing.
+Specifically, it employs a sparse and frozen down-projection matrix AA to compute routing scores $\mathbf{y}=A\mathbf{x}+\mathbf{d}$, where $\mathbf{d}\in\mathbb{R}^{r}$ is a learnable bias term for load balancing.
 Based on these scores, a top-$k$ operator selects the active indices $\mathcal{I}_{\text{top}k}\subseteq\{1,\dots,r\}$.
@@
 As illustrated in Fig.
-4, all Zipper-LoRA variants leverage a globally shared down-projection matrix $A$, ensuring common acoustic-linguistic feature extraction.
+4, all Zipper-LoRA variants leverage a globally shared down-projection matrix AA, ensuring common acoustic-linguistic feature extraction.
 The fundamental distinction between these variants lies in the rank-wise construction of the up-projection matrix $B$.
@@
 This integration can be realized through either a static allocation or a dynamic rank-wise composition.
-To provide a comprehensive understanding of the proposed framework, in the following sections, we first establish the definition of the shared down-projection matrix $A$ and the rank-wise construction of $B$.
+To provide a comprehensive understanding of the proposed framework, in the following sections, we first establish the definition of the shared down-projection matrix AA and the rank-wise construction of $B$.
 We then present the proposed Zipper-LoRA family, ranging from the static language-wise hard split to dynamic binary selection and soft mixing mechanisms.
 This is followed by the LID-aware contextual routing that manages these rank-wise compositions, concluding with the specialized training strategies tailored for optimizing Zipper-LoRA, such as segment-wise encoding and parameter initialization.
-IV-A Definition: Shared $A$ and Rank-wise Construction of $B$
+IV-A Definition: Shared AA and Rank-wise Construction of $B$
 Let $A\in\mathbb{R}^{r\times d_{in}}$ be the shared down-projection matrix.
@@
 We use Whisper Large-v3 [37] as the speech encoder and Qwen3-1.7B [50] as the decoder-only language model.
-The modality projector adopts a gated projection module that performs temporal downsampling by a factor of $4$.
-Specifically, it applies two 1D convolution layers (kernel size $4$, stride $4$) to produce a gate branch and an up-projection branch, fuses them via a SiLU-gated interaction [11], and then applies a linear transformation followed by a residual connection with layer normalization.
+The modality projector adopts a gated projection module that performs temporal downsampling by a factor of 44.
+Specifically, it applies two 1D convolution layers (kernel size 44, stride 44) to produce a gate branch and an up-projection branch, fuses them via a SiLU-gated interaction [11], and then applies a linear transformation followed by a residual connection with layer normalization.
 A final linear projection outputs the speech-conditioned embedding sequence.
@@
 We set the rank to $32$ and the scaling factor to $64$ for all methods.
-For FlyLoRA [57], we activate $8$ rank components.
-For the chunked encoder regime, we randomly sample chunk durations from $\{1,2,4,8\}$ seconds during training and use a fixed chunk duration of $8$ seconds during decoding; no overlap is used.
+For FlyLoRA [57], we activate 88 rank components.
+For the chunked encoder regime, we randomly sample chunk durations from $\{1,2,4,8\}$ seconds during training and use a fixed chunk duration of 88 seconds during decoding; no overlap is used.
 The non-chunked regime encodes each utterance with full context.
-All experiments are trained on $4$ NVIDIA A800 GPUs (80 GB) with DeepSpeed ZeRO stage $2$ [38], using a learning rate of $2\times 10^{-5}$, a warmup ratio of $10\%$, and a cosine learning-rate schedule.
+All experiments are trained on 44 NVIDIA A800 GPUs (80 GB) with DeepSpeed ZeRO stage 22 [38], using a learning rate of $2\times 10^{-5}$, a warmup ratio of $10\%$, and a cosine learning-rate schedule.
 To improve memory efficiency, we pack training samples by constraining the total sequence length to at most 2048 tokens, where the packed length includes the downsampled speech tokens, the text tokens, and reserved prompt tokens; packing is performed within the same language.